# Diffusive and hydrodynamic magnetotransport around a density perturbation in a two-dimensional electron gas

### v1

@@ Abstract
We study current flow around a circular density depletion in a two-dimensional electron gas in the presence of a strong magnetic field. The depletion is parametrized by a power-law tail with an exponent $\beta>2$. We show that current and electrochemical potential are exponentially suppressed inside a surrounding area much larger than the geometric size of the depletion region. The corresponding “no-go” radius grows as a certain power of the magnetic field. Residual current and potential exhibit spiraling patterns inside the no-go region. Outside of it, they acquire corrections inversely proportional to the distance, which is known as the Landauer resistivity dipole. The Landauer dipole is rotated by the angle $\pi(1-1/\beta)$ with respect to the direction of the average electric field. We also consider the effect of electron viscosity and show that the variation of the no-go radius with magnetic field becomes more rapid if viscosity is large enough. In that regime the size of the Landauer dipole is set by the Gurzhi length, which is much larger than the no-go radius, which is in turn much larger than the geometric size of the depletion. Our results may be useful for interpreting nanoimaging of current distribution in graphene and other two-dimensional systems.
@@ I Introduction
Direct imaging of electron transport in two-dimensional (2D) systems has become possible recently with the development of local probes capable of nanometer-scale resolution. Scanning single-electron transistors can measure local compressibility and potential [1, 2, 3, 4]. Subsurface charge accumulation imaging and scanning capacitance microscopy have been used to image local potentials and dissipation [5, 6, 7]. Scanning atomic force microscopy [8], scanning tunneling microscopy and potentiometry [9, 10, 11], scanning SQUID [12, 13], and nitrogen-vacancy center magnetometry [14] are also among the techniques that provide access to the distributions of the electrochemical potential $e\Phi$ and current density $\mathbf{j}$. Strong magnetic field has been shown to have a dramatic effect on electron transport. This regime can be defined by the condition that the Hall conductivity $\sigma_{xy}$ exceeds the dissipative conductivity $\sigma_{xx}$. In particular, in the quantum Hall effect (QHE) the latter vanishes, and as a result, current can flow only along the edges. In real QHE devices, the longitudinal conductivity $\sigma_{xx}$ is small but finite, so the current penetrates into the bulk [15, 16]. The current distribution is nontrivial and highly sensitive to device geometry and inhomogeneities which can be modeled by spatially varying $\sigma_{xx}$ and $\sigma_{xy}$ [17].
@@
$\Phi(r,\theta)\propto-x-\frac{R_{L}^{2}}{r}\cos\left(\theta+\theta_{L}\right),$ (1)
where the first term corresponds to a uniform electric field in the $\hat{\mathbf{x}}-$direction and the second is a correction known as the Landauer resistivity dipole [22, 23, 24, 25, 26]. (A 2D dipole potential behaves as $1/r$ rather than $1/r^{2}$ in three dimensions.) In the absence of the magnetic field, $B=0$, the dipole had the angular orientation $\theta_{L}=0$, consistent with a simple qualitative picture where the current incident on the obstacle causes charge crowding (uncrowding) on its upstream (downstream) sides. In the presence of a weak magnetic field, $B\lesssim 1\,\mathrm{T}$, the dipole rotated with respect to the asymptotic direction of the electric field: $\theta_{L}\neq 0$. This implies a twisted, spiral-like current flow around the depletion, Fig. 1(a). Additionally, the region where the current was suppressed grew beyond the geometric boundary of the ppnn junction and acquired some spiral-like structure. These observations were analyzed using several theoretical approaches: an analytical model of diffusive transport with the hard-wall boundary condition at the ppnn junction, a semiclassical ballistic correction to this model, and numerical simulations of quantum transport. The first one predicted that the Landauer dipole rotation angle is $\theta_{L}=2\theta_{H}$ where $\theta_{H}=\arctan(\sigma_{xy}/\sigma_{xx})$ is the Hall angle, in agreement with the experiment. The last two approaches explained an additional feature of $\Phi(x,y)$ detected near the obstacle in strong fields $B\gtrsim 1\,\mathrm{T}$: a small kink at the distance of one cyclotron diameter away from the ppnn boundary.
@@
If the density depletion is induced by local doping, then according to electrostatics [27], $\beta=3$ at large $r$. Other values of this exponent may be possible to realize using electrostatic doping with gate dielectrics of varying thickness or electrodes of special shape. We will assume that the power-law decrease is fast enough, $\beta>2$. Under this condition the depletion acts as a local perturbation so that the asymptotic dipolar law [Eq. (1)] is valid.
Our analysis shows that the size $R_{L}$ and the rotation angle $\theta_{L}$ of the Landauer resistivity dipole are qualitatively different in the step-like and the gradual density profile cases. In the latter, $R_{L}$ is much larger than $a$ and grows with $B$ if the magnetic field is strong. This implies that even a weak density depletion can strongly repel the injected current. The size of the “no-go” region, i.e., the region where the current is strongly suppressed compared to its asymptotic value at infinity is of the order of $R_{\mathrm{dip}}$ rather than $a$ and it grows with $B$. Remarkably, the Landauer dipole rotation angle is reduced compared to that for the hard-wall case: $\theta_{L}\simeq 2\theta_{H}-(\pi/\beta)$.
It has been shown theoretically [28, 29, 30] that the size of the Landauer dipole is determined primarily by the so-called Gurzhi length if the electron fluid has high viscosity. This effect is related to the Stokes paradox in the context of a hydrodynamic flow around a cylinder [31, 32]. High-viscosity electron fluid can be realized in modern high-mobility 2D systems, such as semiconductor nanostructures [33, 34, 35, 36, 37, 38, 39, 40] and graphene [41, 42, 43, 44, 45, 14, 46, 47], and also in some quasi-2D metals [48, 49, 50, 51, 52] where electron-electron scattering rate is much higher than electron-impurity one. Motivated by this, we examine the role of viscosity in our magnetotransport problem. We show that in the viscosity-dominated regime, the scaling of the no-go radius as a function of $B$ is modified compared to the diffusion-dominated one: it is characterized by a significantly larger power-law exponent. The size of the Landauer dipole is given by the Gurzhi length, with a logarithmic correction, similar to the zero-field case.
@@ II.3 Analytical solutions
cf. Eqs. (1) and (19). Substituting Eq. (17) into Eq. (9), we find
@@
In addition to Eq. (18), we need a boundary condition for short distances. For example, in the model where the conductivity vanishes at the origin [Eq. (13)], the suitable boundary condition is $\phi(0)=0$. In the model with a finite-size insulating core, i.e., $n(r)=\sigma_{xx}(r)=\sigma_{xy}(r)=0$ at $r\leq a$, the potential $\phi(a)$ does not have to vanish but must be bounded. (In fact, it is exponentially small if $\alpha\ll 1$, see below.) To fix $\phi(r)$ at $r<a$ one needs to make additional assumptions because Eq. (20) becomes an identity at such $r$. For simplicity, we choose the linear ansatz,
$\phi(r)=\frac{r}{a}\,\phi(a),\quad r<a\,$, (21)
which corresponds to a uniform electric field of magnitude $|\phi(a)/a|$.
@@
A closer look at the no-go region reveals that $\phi(r)$ has some oscillations. The period of the oscillations becomes shorter and their amplitude becomes exponentially small as $r$ decreases. This fine structure is put on view more clearly in a false color plot of the 2D potential distribution $\Phi(x,y)$, Fig. 3. The superimposed equipotential contours are seen to spiral inward clockwise as they approach the insulating core from infinity and then spiral outward counterclockwise as they go away to infinity. (Inside the $r<a$ core, the equipotentials go across as straight lines but this is not shown in Fig. 3.) This behavior is consistent with the qualitative picture introduced in Sec. II.2. Solutions exhibiting similar spiral patterns have also been discussed in Ref. [53].
@@
which can be interpreted as a radial Schrödinger equation in the $p$-wave channel for a particle subject to an imaginary potential $V(r)\propto-ir^{-\beta-2}$. This equation has two linearly independent solutions, one of which exponentially increases and one decreases as $r$ approaches the origin:
@@
$\theta_{L}=\pi-2\arctan\alpha\simeq\pi\,.$ (27)
Comparing Eqs. (25) and (27) we see that for the gradual density profile the angle $\theta_{L}$ is reduced by $\pi/\beta$ compared to that for the hard-wall barrier.
Note that the hard-wall case can be analyzed within our formalism simply by taking the limit $\beta\to\infty$ where the density profile becomes step-like: $n(r)=0$ at $r<a$ and $n(r)=n_{0}$ at $r>a$. We have checked that Eq. (27) is then fully reproduced provided we use the exact solutions of Eq. (20) rather than the small-$\alpha$ approximation of Eq. (24), see Appendix A.
@@ III.1 Model
where $v_{F}$ is the Fermi velocity and $\tau_{\mathrm{ee}}$ is the average time between collisions. (For the specific case of graphene, see also Refs. 58, 59, 60, 61.) We focus on the range of magnetic fields such that $1/\tau_{\mathrm{mr}}\ll\omega_{c}\ll 1/\tau_{\mathrm{ee}}$ where $\nu$ is $B$-independent. The corresponding hierarchy of the key length scales is
@@
where $\mathbf{u}$ is the flow velocity related to the current via
@@
and $\nu_{B}$ is the bulk viscosity. To keep the model as simple as possible, we assume that $\nu_{B}+\frac{1}{2}\nu=0$, so that the total viscous friction force is just
@@
In principle, our system also possesses Hall viscosity $\nu_{H}=2\nu\omega_{c}\tau_{\mathrm{ee}}$ [54, 55, 56, 57, 62, 63, 64]. Under the stated condition $\omega_{c}\tau_{\mathrm{ee}}\ll 1$ we can neglect it. Our Eq. (30) [and the earlier Eq. (3) as well] also neglects the thermoelectric effect arising from the gradient of temperature $T$ that enters the Gibbs-Duhem relation $\bm{\nabla}\varepsilon_{F}=n\bm{\nabla}p-s\bm{\nabla}T$. Here $s$ is the entropy per particle and $p$ is the pressure. This effect should be small in a degenerate 2D electron fluid to begin with and it results in corrections that are of higher order in gradients than the terms we have kept, cf. Refs. [65, 60, 61, 29]. On the other hand, Eq. (30) does include the term $\bm{\nabla}p$ appearing in the hydrodynamical equation of motion [31, 32] because $\mathbf{E}$ is not the electric field but rather the electrochemical gradient, see Sec. II.
@@
is referred to as the Gurzhi length [66]. As we discuss below, viscosity plays an important role if $l_{G}\gg a$.
@@ III.3 Far-field region
$\Psi(\mathbf{r})=\Re\mathrm{e}\left[\psi(r)ie^{i(\theta+\theta_{H})}\right],$ (41)
@@
The only effect of the magnetic field is to rotate the Landauer dipole by twice the Hall angle $\theta_{H}$ with respect to the average electric field,
@@
same as for diffusive transport around a hard-wall obstacle. (However, if the Hall viscosity $\nu_{H}$ is retained and the boundary conditions are different from the no-slip ones [67], then there is an additional correction to $\theta_{L}$ of the order of $\nu_{H}/\nu\sim\omega_{c}\tau_{\mathrm{ee}}$ [29, 30].)
If $l_{G}$ is much larger than $a$, we can expand the Bessel functions in Eq. (44) to get

### v2

@@ Abstract
We study current flow around a density inhomogeneity in a two-dimensional electron gas in the presence of a strong magnetic field. The inhomogeneity is parametrized by a power-law tail with an exponent $\beta>2$. We show that current and electrochemical potential are exponentially suppressed inside a surrounding area much larger than the geometric size of the perturbation. The corresponding “no-go” radius grows as a certain power of the magnetic field. Residual current and potential exhibit spiraling patterns inside the no-go region. Outside of it, they acquire corrections inversely proportional to the distance, which is known as the Landauer resistivity dipole. The Landauer dipole is rotated by the angle $\pi(1-1/\beta)$ with respect to the average electric field. The rotation direction depends on whether the local density is raised or lowered. We also consider the effect of electron viscosity and show that the variation of the no-go radius with magnetic field becomes more rapid if viscosity is large enough. The Landauer dipole size is set by the Gurzhi length, which is much larger than the no-go radius, which is in turn much larger than the geometric size of the perturbation. Our results may be useful for interpreting nanoimaging of current distribution in graphene and other two-dimensional systems.
@@ I Introduction
Direct imaging of electron transport in two-dimensional (2D) systems has become possible recently with the development of local probes capable of nanometer-scale resolution. Scanning single-electron transistors can measure local compressibility and potential [1, 2, 3, 4]. Subsurface charge accumulation imaging and scanning capacitance microscopy have been used to image local potentials and dissipation [5, 6, 7]. Scanning atomic force microscopy [8], scanning tunneling microscopy and potentiometry [9, 10, 11], scanning SQUID [12, 13], and nitrogen-vacancy center magnetometry [14, 15] are also among the techniques that provide access to the distributions of the electrochemical potential $e\Phi$ and current density $\mathbf{j}$. Strong magnetic field has been shown to have a dramatic effect on electron transport. This regime can be defined by the condition that the Hall conductivity $\sigma_{xy}$ exceeds the dissipative conductivity $\sigma_{xx}$. In particular, in the quantum Hall effect (QHE) the latter vanishes, and as a result, current can flow only along the edges. In real QHE devices, the longitudinal conductivity $\sigma_{xx}$ is small but finite, so the current penetrates into the bulk [16, 17]. The current distribution is nontrivial and highly sensitive to device geometry and inhomogeneities which can be modeled by spatially varying $\sigma_{xx}$ and $\sigma_{xy}$ [18].
@@
$\Phi(r,\theta)\propto-r\cos\theta-\frac{R_{L}^{2}}{r}\cos\left(\theta+\theta_{L}\right),$ (1)
where the first term corresponds to a uniform electric field in the $\hat{\mathbf{x}}-$direction and the second is a correction known as the Landauer resistivity dipole [23, 24, 25, 26, 27, 28, 29]. (A 2D dipole potential behaves as $1/r$ rather than $1/r^{2}$ in three dimensions. Here $r=|\mathbf{r}|$ is the radial distance and $\theta$ is the polar angle.) In the absence of the magnetic field, $B=0$, the dipole had the angular orientation $\theta_{L}=0$, consistent with a simple qualitative picture where the current incident on the obstacle causes charge crowding (uncrowding) on its upstream (downstream) sides. In the presence of a weak magnetic field, $B\lesssim 1\,\mathrm{T}$, the dipole rotated with respect to the asymptotic direction of the electric field: $\theta_{L}\neq 0$. This implies a twisted, spiral-like current flow around the depletion, Fig. 1(a). Additionally, the region where the current was suppressed grew beyond the geometric boundary of the ppnn junction and acquired some spiral-like structure. These observations were analyzed using several theoretical approaches: an analytical model of diffusive transport with the hard-wall boundary condition at the ppnn junction, a semiclassical ballistic correction to this model, and numerical simulations of quantum transport. The first one predicted that the Landauer dipole rotation angle is $\theta_{L}=2\theta_{H}$ where $\theta_{H}=\arctan(\sigma_{xy}/\sigma_{xx})$ is the Hall angle, in agreement with the experiment (see also Refs. [25, 28]). The last two approaches explained an additional feature of $\Phi(x,y)$ detected near the obstacle in strong fields $B\gtrsim 1\,\mathrm{T}$: a small kink at the distance of one cyclotron diameter away from the ppnn boundary.
@@
If the density depletion is induced by local doping, then according to electrostatics [30], $\beta=3$ at large $r$. Other values of $\beta$ may be possible to realize using electrostatic doping with gate dielectrics of varying thickness or electrodes of special shape. We will assume that the power-law decrease is fast enough, $\beta>2$. Under this condition, the depletion acts as a local perturbation so that the asymptotic dipolar law [Eq. (1)] is valid.
Our analysis shows that the size $R_{L}$ and the rotation angle $\theta_{L}$ of the Landauer resistivity dipole are qualitatively different in the step-like and the gradual density profile cases. In the latter, $R_{L}$ is much larger than $a$ and grows with $B$ if the magnetic field is strong. This implies that even a weak density depletion can strongly repel the injected current. The size of the “no-go” region, i.e., the region where the current is strongly suppressed compared to its asymptotic value at infinity is of the order of $R_{L}$ rather than $a$ and it grows with $B$. Remarkably, the Landauer dipole rotation angle is reduced compared to that for the hard-wall case: $\theta_{L}\simeq 2\theta_{H}-(\pi/\beta)$. We also briefly examine the case of density accumulation, where the sign of the correction in Eq. (2) is positive, in which case the Landauer dipole rotates in the opposite direction.
It has been shown theoretically [31] that the size of the Landauer dipole is determined primarily by the so-called Gurzhi length [32] if the electron fluid has high viscosity. This effect is related to the Stokes paradox in the context of a hydrodynamic flow around a cylinder [33, 34]. Hydrodynamic flow around a hard obstacle in a 2D electron gas in the presence of a magnetic field has been studied in Refs. [35, 36]. In particular, the charge density distribution of the induced Landauer dipole was investigated in Ref. [35]. High-viscosity electron fluid can be realized in modern high-mobility 2D systems, such as semiconductor nanostructures [37, 38, 39, 40, 41, 42, 43, 44] and graphene [45, 46, 47, 48, 49, 14, 50, 51], and also in some quasi-2D metals [52, 53, 54, 55, 56] where electron-electron scattering rate is much higher than electron-impurity one. Motivated by these developments, we examine the role of viscosity in our magnetotransport problem where the electron density is nonuniform, Eq. (2). We show that in the viscosity-dominated regime, the scaling of the no-go radius as a function of $B$ is modified compared to the diffusion-dominated one: it is characterized by a significantly larger power-law exponent. The size of the Landauer dipole is given by the Gurzhi length, with a logarithmic correction, similar to the zero-field case.
@@ II.2 Qualitative discussion
Note that the advection-diffusion analogy is a fairly faithful representation of the actual physics of the problem. The fictitious flow velocity $\mathbf{v}$ is proportional to the drift velocity $c\left(\mathbf{E}_{0}\times\hat{\mathbf{z}}\right)/B$ of the guiding center of an electron’s cyclotron orbit in crossed electric and magnetic fields. Here $\mathbf{E}_{0}=\bm{\nabla}\varepsilon_{F}/e$ is the electric field that exists in the depletion region in equilibrium.
@@ II.3 Analytical solutions
cf. Eq. (1). Substituting Eq. (17) into Eq. (9), we find
@@
In addition to Eq. (18), we need a boundary condition for short distances. For example, in the model where the conductivity vanishes at a single point [Eq. (13)], the suitable boundary condition is $\phi(0)=0$. In the model with a finite-size insulating core, i.e., $n(r)=\sigma_{xx}(r)=\sigma_{xy}(r)=0$ at $r\leq a$, the potential $\phi(a)$ does not have to vanish but must be bounded. (In fact, it is exponentially small if $\alpha\ll 1$, see below.) To fix $\phi(r)$ at $r<a$ one needs to make additional assumptions because Eq. (9) becomes an identity at such $r$. For example, if electrons are strictly forbidden to enter the core, as in the antidot problem, then $\phi(r)$ would exhibit an inverse square-root divergence (rounded on the scale of the screening length) as $r$ approaches $a$ from inside [58]. To avoid such complications, we assume that the region $r<a$ has a very small uniform conductivity [see Eq. (29) below]. In this case the electric field at $r<a$ is also uniform and the potential is linear,
$\phi(r)=\frac{r}{a}\,\phi(a),\quad r<a\,.$ (21)
@@
A closer look at the no-go region reveals that $\phi(r)$ has some oscillations. The period of the oscillations becomes shorter and their amplitude becomes exponentially smaller as $r$ decreases. This fine structure is put on view more clearly in a false color plot of the 2D potential distribution $\Phi(x,y)$, Fig. 3. The superimposed equipotential contours are seen to spiral inward clockwise as they approach the insulating core from $y=+\infty$ and then spiral outward counterclockwise as they go away to $y=$-$\infty$. (Inside the $r<a$ core, the equipotentials go across as straight lines but this is not shown in Fig. 3.) This behavior is consistent with the qualitative picture introduced in Sec. II.2. Solutions exhibiting similar spiral patterns have also been discussed in Ref. [57].
@@
which can be interpreted as a radial Schrödinger equation in the $p$-wave channel for a particle subject to an imaginary scalar potential $V(r)\propto-ir^{-\beta-2}$. This equation has two linearly independent solutions, one of which exponentially increases and one decreases as $r$ approaches the origin:
@@
$\theta_{L}=2\theta_{H}=2\arctan\left(\frac{1}{\alpha}\right)\simeq\pi\,$, (27)
see also Eq. (31) below. Comparing Eqs. (25) and (27), we discover that for the gradual density profile the angle $\theta_{L}$ is reduced by $\pi/\beta$ compared to that for the hard-wall barrier.
Note that the hard-wall case can be analyzed within our formalism simply by taking the limit $\beta\to\infty$ where the density profile becomes step-like: $n(r)=0$ at $r<a$ and $n(r)=n_{0}$ at $r>a$. Using the exact solutions of Eq. (20) rather than the small-$\alpha$ approximation of Eq. (24), we can reproduce the formula $\theta_{L}=2\theta_{H}$ in full, see Appendix A.
II.4 Density accumulation
Let us briefly discuss the model where the electron density is locally raised rather than depleted,
$\frac{n(r)}{n_{0}}\simeq 1+\left(\frac{a}{r}\right)^{\beta},\quad r\gg a.$ (28)
In this case the direction of the drift velocity $\mathbf{v}=\hat{\mathbf{z}}\times\bm{\nabla}\sigma_{xy}(\mathbf{r})$ is reversed. The corresponding plot of the electrochemical potential can be obtained by reflecting Fig. 3 across the $x$-axis. Equations (22)–(24) are replaced by their complex conjugates and the sign of $\theta_{L}$ in Eq. (25) is reversed.
As for the hard-wall limit, the closest related model is an abrupt junction:
[…]
Matching the outer and inner solutions, Eqs. (18) and (21), at the boundary $r=a$, we find
$\lambda=\frac{\left(\sigma_{xx}^{(0)}+i\sigma_{xy}^{(0)}\right)-\left(\sigma_{xx}^{(1)}+i\sigma_{xy}^{(1)}\right)}{\left(\sigma_{xx}^{(0)}-i\sigma_{xy}^{(0)}\right)+\left(\sigma_{xx}^{(1)}+i\sigma_{xy}^{(1)}\right)}\,a^{2}\,.$ (30)
If $\sigma_{xx}^{(0)}/\sigma_{xy}^{(0)}=\sigma_{xx}^{(1)}/\sigma_{xy}^{(1)}=\alpha$, then
$\lambda=-\frac{1-i\alpha}{1-iA\alpha}\,a^{2}\,,\quad A\equiv\frac{\sigma_{xx}^{(1)}+\sigma_{xx}^{(0)}}{\sigma_{xx}^{(1)}-\sigma_{xx}^{(0)}}\,.$ (31)
For example, if the density accumulation is very high, $A\simeq 1$, then the Landauer dipole size $R_{L}\simeq a$ and angle $\theta_{L}\simeq-\pi$ are approximately independent of the magnetic field. For more details, see Appendix A.
@@ III.1 Model
where $v_{F}$ is the Fermi velocity and $\tau_{\mathrm{ee}}$ is the average time between collisions. (For the specific case of graphene, see Refs. 63, 64, 65, 66, 67.) We focus on the range of magnetic fields such that $1/\tau_{\mathrm{mr}}\ll\omega_{c}\ll 1/\tau_{\mathrm{ee}}$ where $\nu$ is $B$-independent. The corresponding hierarchy of the key length scales is
@@
where $d=2$ is the space dimension, $\mathbf{u}$ is the flow velocity related to the current via
@@
and $\nu_{b}$ is the bulk viscosity (or “second” viscosity [34]). For simplicity, we assume that $\nu_{b}+\frac{1}{2}\nu=0$, so that the total viscous friction force is just
@@
In principle, the system also possesses the Hall viscosity $\nu_{H}=2\nu\omega_{c}\tau_{\mathrm{ee}}$ [59, 60, 61, 62, 68, 69, 70]. Under the stated condition $\omega_{c}\tau_{\mathrm{ee}}\ll 1$ we can neglect it. Our Eq. (34) and the earlier Eqs. (3), (7) also neglect the thermopower arising from the gradient of temperature $T$ that enters the Gibbs-Duhem relation $\bm{\nabla}\varepsilon_{F}=n^{-1}\bm{\nabla}p-s\bm{\nabla}T$. Here $s$ is the entropy per particle and $p$ is the pressure. Thermopower may lead to experimentally observable effects in the low-density (or $p$-$n$ junction) region $r\sim a$ [21, 22]; otherwise, it is rather small in a degenerate 2D electron fluid and it results in corrections that are of higher order in gradients than the terms we have kept, cf. Refs. [71, 65, 66, 67]. On the other hand, Eq. (34) does include the pressure contribution $p\delta_{ik}$ to the total stress tensor [33, 34] because $\mathbf{E}=-\bm{\nabla}\Phi_{e}-\frac{1}{e}\,\bm{\nabla}\varepsilon_{F}$ is not the electric field but rather the electrochemical gradient.
@@
is the aforementioned Gurzhi length. As we discuss below, viscosity plays an important role if $l_{G}\gg a$.
@@ III.3 Far-field region
$\Psi(\mathbf{r})=\Re\mathrm{e}\left[-ie^{i(\theta+\theta_{H})}\psi(r)\right],$ (45)
@@
Similar formulas have been derived [35, 67, 36] for other types of boundary conditions, e.g., no-stress boundary conditions. The only effect of the magnetic field is to rotate the Landauer dipole by twice the Hall angle $\theta_{H}$ with respect to the average electric field,
@@
same as for diffusive transport around a hard obstacle. (If the Hall viscosity $\nu_{H}$ is retained, then the flow-induced density profile is rotated by an angle of the order of $\nu_{H}/\nu\sim\omega_{c}\tau_{\mathrm{ee}}$ [35].) If $l_{G}$ is much larger than $a$, we can expand the Bessel functions in Eq. (48) to get